Brown Co. purchased a piece of equipment at the beginning of Year 1 for $25,000,000. Management estimates that the equipment will have a useful life of seven years and a $5,000,000 salvage value. The depreciation expense recorded for tax purposes is computed using the double-declining balance method of depreciation. The company uses the straight-line method of depreciation for reporting purposes.

Calculate the amount of depreciation expense for reporting purposes for Year 4 (i.e., the fourth full year of depreciation). Then calculate the amount of depreciation expense for tax purposes for Year 4 (i.e., the fourth full year of depreciation). Use this information to determine the correct answers to the following two (2) questions:

2.) Assuming that the company’s tax rate is 25%, what dollar amount (as an absolute value) will be added (if your answer to the prior question was “a”) or subtracted (if your answer to the prior question was “b”) to the deferred tax account in Year 4 as a result of the depreciation timing difference?
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Thanks!
Brown Co Straight Line Table Cost/Opening Book Value Depreciation Expense Accumulated depreciation Net Book Value
Cost of Equipment    25,000,000.00 A Year 0                        25,000,000.00                                      -                                                    -             25,000,000.00
Residual Value      5,000,000.00 B Year 1                        25,000,000.00                  2,857,142.86                              2,857,142.86           22,142,857.14
Depreciable Value 20,000,000.00 C=A-B Year 2                        22,142,857.14                  2,857,142.86                              5,714,285.71           19,285,714.29
Year 3                        19,285,714.29                  2,857,142.86                              8,571,428.57           16,428,571.43
Straight Line Method Year 4                        16,428,571.43                  2,857,142.86                            11,428,571.43           13,571,428.57
Depreciable Value    20,000,000.00 See C Total               11,428,571.43
Life                      7.00 D
Annual depreciation      2,857,142.86 E=C/D
M N=M*L
Double Declining Method Table Cost/Opening Book Value Depreciation Expense Accumulated depreciation Net Book Value
Cost of Equipment    25,000,000.00 See A Year 0                        25,000,000.00                                      -                                                    -             25,000,000.00
Life                      7.00 See D Year 1                        25,000,000.00                  7,142,857.14                              7,142,857.14           17,857,142.86
Annual depreciation      3,571,428.57 J=A/D Year 2                        17,857,142.86                  5,102,040.82                            12,244,897.96           12,755,102.04
Depreciation rate 14.29% K=J/A Year 3                        12,755,102.04                  3,644,314.87                            15,889,212.83             9,110,787.17
Double Depreciation % 28.57% L=K*2 Year 4                          9,110,787.17                  2,603,082.05                            18,492,294.88             6,507,705.12
Total               18,492,294.88
Depreciation for 4th year under Straight Line Method            2,857,142.86
Depreciation for 4th year under Double Declining Method            2,603,082.05
Depreciation charged less under taxation               254,060.81
It means income is more by               254,060.81
Tax Rate 25%
Extra tax paid                 63,515.20
This is differed tax asset.
So $ 63,515.20 will be added to the to the deferred tax account in Year 4.